# From Memories to Maps: Mechanisms of In-Context Reinforcement Learning in Transformers

### v2

@@ A.2 Tree Mazes
Appendix B Model and Training Parameters
We largely follow the same architecture as that of Lee et al. [2023], a GPT-2 style model with causal attention and without positional embeddings. Our default model has 4 heads, 3 layers, and embedding dimension of $512$. Context memory tokens consist of the $(s_{t},a,s^{\prime}_{t},r_{t})$ tuple concatenated together into one vector. Thus, tokens are $26$-dimensional in gridworld and $25$-dimensional in tree maze. The query token consists of $(s_{q},\vec{0})$ for query state $s_{q}$, where $\vec{0}$ provides 0-padding to reach the desired vector size. These tokens are projected into model embedding space through a learnable linear layer. The model samples greedily in the gridworld environment and with softmax sampling in the tree maze environment (both settings were empirically determined).
In contrast to Lee et al. [2023], we provide the query token at the end of the context memory. This is to allow a clearer interpretation in which context memory tokens represent previous experiences in the maze that are stored in episodic memory. The query token is only provided at decision time, and the agent must use previous memories to guide its present decision.
To allow for query tokens at the end of an input sequence sequence and to preserve efficient pretraining, we make modifications to the pretraining procedure of Lee et al. [2023], which we describe here. Let’s say we have a pretraining task with context memory tokens $\mathcal{D}$ and query state $s_{q}$. To encourage length generalization, we would like to train the model on many sequence lengths– let’s say every $t_{step}$ timesteps of $\mathcal{D}$. To do so from one forward pass, we first construct a sequence $\mathcal{D}_{train}$ where $s_{q}$ is interleaved every $t_{step}$ timesteps of $\mathcal{D}$: $\mathcal{D}_{train}=[\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{t_{step}},s_{q},\mathcal{D}_{t_{step}+1},\dots\mathcal{D}_{2*t_{step}},s_{q},\mathcal{D}_{2*t_{step}+1},\dots,\mathcal{D}_{T},s_{q}]$. We then construct an attention mask $A_{mask}=A_{causal}+A_{query}$, where $A_{causal}$ is the standard causal attention mask with $\infty$- ∞ values in the upper-triangular and 00 elsewhere. $A_{query}$ ensures that query tokens are not processed by context memory tokens by masking columns corresponding to the query token:
@@ Appendix B Model and Training Parameters
We use Adam optimizer with weight decay $1\times 10^{-5}$. We use a batch size of $1024$ for gridworld and a batch size of $512$ for tree mazes. We train the model for 25 epochs in gridworld and 50 epochs in tree mazes. We use a learning rate of $1\times 10^{-4}$ which we linearly decrease to $1\times 10^{-5}$ over the course of training (we found that this empirically worked well). For each training run, we use two NVIDIA H100 GPUs. This results in around $1$ hour of training time for gridworld and $1.5$ hours of training time for tree mazes. For each task, we train models from $5$ random initializations, for both a dropout of 00 and $0.2$. We then select the best model via validation performance for each task. This selected model is the model we analyze for each task.
@@
Appendix C Results Sensitivity to Task/Model Parameters
@@
Appendix D Additional Behavioral Results in Gridworld and Tree Mazes
Here, we show additional learning results for both the gridworld task and tree mazes. We first show additional in-context learning curves for gridworld (Fig 12). We note that in-context learning can still be unstable at times. In part, this may be because the model is sometimes tested on states it has not experienced. Thus, it is more difficult to navigate into previously experienced territory to find the goal. We also suspect that improvements in the training procedure or architecture that we have not explored could also produce a more performant model.
@@ Appendix D Additional Behavioral Results in Gridworld and Tree Mazes
[…]
@@
Appendix E Q-Learning Simulations
@@
There are also a few additional training details for the DQN. We randomly reinitialize the network weights at each context length before we run the training procedure. This is because we find that resetting the weights works empirically better than initializing with the weights from the previous context length the model was trained on (this is reasonable, as the latter induces a continual learning problem). To maintain as many parallels to standard methods as possible, we also adopt a double deep Q learning framework [Van Hasselt et al., 2016]. We use a target network that is updated every $10$ epochs. We don’t think this detail is critical (and empirically the use of a target network here doesn’t seem to impact performance) as the learning problem in our setting is fully stationary.
@@
Appendix F Additional in-context representation learning results
@@
Appendix G Kernel Alignment
@@
Appendix H Analyzing Cross-Context Learning in Gridworld
@@
Appendix I Linear Decoder Setup
@@
To fit a linear decoder, we use ridge regression. We standardize features to 00 mean and unit variance. The regularization strength $\alpha$ was selected through 5-fold cross-validation using a grid search over regularization strengths from $[10^{0},10^{4}]$, with 10 logarithmically-spaced values. Cross-validation was performed with shuffled splits. For each $\alpha$, we computed the mean $R^{2}$ score across all validation folds and selected the $\alpha$ that maximized this cross-validation performance. The final decoder for each layer was fit on the complete training set using the $\alpha$ found previously. Model performance was evaluated on the held-out test set.
For circular variables such as angles, we cannot directly apply standard regression since the circular nature of the data violates the assumptions of linear models (e.g., an angle of $\pi$ and $-\pi$ represent the same direction but appear numerically distant). Instead, we decompose each target angle $\theta$ into its sine and cosine components: $\text{sin}(\theta)$ and $\text{cos}(\theta)$. We then fit two separate ridge regressors to predict these components independently, using the same cross-validation procedure described above. To obtain the final angle prediction, we convert the predicted sine and cosine values back to angles using the arctangent function: $\hat{\theta}=\text{arctan2}(sin(\hat{\theta}),cos(\hat{\theta}))$.
@@
Appendix J Gradient Attribution Method
To get gradient attributions, we use integrated gradients [Sundararajan et al., 2017]. As a reminder, the model output is a vector defining weights over actions. We calculate the gradient of the model’s output for the optimal action with respect to input tokens. We define the baseline inputs as the original context memory dataset $\mathcal{D}$ but with actions ablated (that is, $a=\textbf{0}$). We integrate over $20$ steps.
Appendix K Tests for Model-Free Reasoning
@@
Appendix L Tests for Model-Based Reasoning
@@
We next probe for signatures of model-based reasoning. That is, we look for evidence that the model utilizes path planning to choose the correct action from the query state. This is connected to questions of state tracking [Li et al., 2025] and understanding how models simulate successive transitions between states. Li et al. [2025] propose different ways that transformer models can do this path planning, from forward rollouts to more sophisticated, mergesort-like algorithms. Each of these algorithms require simulating transitions through intermediate states between query and goal. Thus, to test for the presence of path planning, we look for evidence that information about intermediate states are utilized at decision time. Specifically, we isolate decision time as computations conducted in the query token stream.
@@
Appendix M Gridworld Mechanisms
@@
Here, we show additional decoding results for gridworld. In the main text, we discuss how XY location and angle to goal can be decoded clearly from the query token stream of the model. We also find that these variables can be decoded from the context tokens. We find this interesting as it connects to our findings in tree mazes where memory tokens contain not just information about the original event (i.e., transition), but also additionally computed features.
Appendix N Tree Maze Mechanisms
Here, we show additional results for our tree maze analysis. In the tree maze task, the optimal action is often to transition to the parent node (specifically, this is true in all but the $6$ states that comprise the L-R path). This bias is reflected in the model. Without reward information, the model defaults to transition towards the parent node unless it is at the root node (Fig 23AB). Thus, we believe the model takes its default action unless it accumulates enough evidence through its layer computations to do otherwise. As discussed in the main text, we think this is done by tagging context tokens on the L-R path. The model can then aggregate this information to determine whether or not a query state is on the L-R path (and thus has an optimal action of left or right). The additional sufficiency experiments we conduct as an extension of Fig 7C seems to support this. As long as the query token in the last layer can attend to context tokens that (1) themselves contain the query token and (2) are on the L-R path going towards or away from goal, the model output is unaffected (Fig 23C-F).

### v3

@@ B.2 Tree Mazes
Appendix C Model and Training Parameters
We largely follow the same architecture as that of lee2023supervised, a GPT-2 style model with causal attention and without positional embeddings. Our default model has 4 heads, 3 layers, and embedding dimension of $512$. Context memory tokens consist of the $(s_{t},a,s^{\prime}_{t},r_{t})$ tuple concatenated together into one vector. Thus, tokens are $26$-dimensional in gridworld and $25$-dimensional in tree maze. The query token consists of $(s_{q},\vec{0})$ for query state $s_{q}$, where $\vec{0}$ provides 0-padding to reach the desired vector size. These tokens are projected into model embedding space through a learnable linear layer. The model samples greedily in the gridworld environment and with softmax sampling in the tree maze environment (both settings were empirically determined).
In contrast to lee2023supervised, we provide the query token at the end of the context memory. This is to allow a clearer interpretation in which context memory tokens represent previous experiences in the maze that are stored in episodic memory. The query token is only provided at decision time, and the agent must use previous memories to guide its present decision.
To allow for query tokens at the end of an input sequence sequence and to preserve efficient pretraining, we make modifications to the pretraining procedure of lee2023supervised, which we describe here. Let’s say we have a pretraining task with context memory tokens $\mathcal{D}$ and query state $s_{q}$. To encourage length generalization, we would like to train the model on many sequence lengths– let’s say every $t_{step}$ timesteps of $\mathcal{D}$. To do so from one forward pass, we first construct a sequence $\mathcal{D}_{train}$ where $s_{q}$ is interleaved every $t_{step}$ timesteps of $\mathcal{D}$: $\mathcal{D}_{train}=[\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{t_{step}},s_{q},\mathcal{D}_{t_{step}+1},\dots\mathcal{D}_{2*t_{step}},s_{q},\mathcal{D}_{2*t_{step}+1},\dots,\mathcal{D}_{T},s_{q}]$. We then construct an attention mask $A_{mask}=A_{causal}+A_{query}$, where $A_{causal}$ is the standard causal attention mask with -$\infty$ values in the upper-triangular and 0 elsewhere. $A_{query}$ ensures that query tokens are not processed by context memory tokens by masking columns corresponding to the query token:
@@ Appendix C Model and Training Parameters
We use Adam optimizer with weight decay $1\times 10^{-5}$. We use a batch size of $1024$ for gridworld and a batch size of $512$ for tree mazes. We train the model for 25 epochs in gridworld and 50 epochs in tree mazes. We use a learning rate of $1\times 10^{-4}$ which we linearly decrease to $1\times 10^{-5}$ over the course of training (we found that this empirically worked well). For each training run, we use two NVIDIA H100 GPUs. This results in around $1$ hour of training time for gridworld and $1.5$ hours of training time for tree mazes. For each task, we train models from $5$ random initializations, for both a dropout of 0 and $0.2$. We then select the best model via validation performance for each task. This selected model is the model we analyze for each task.
@@
Appendix D Results Sensitivity to Task/Model Parameters
@@
Appendix E Additional Behavioral Results in Gridworld and Tree Mazes
Here, we show additional learning results for both the gridworld task and tree mazes. We first show additional in-context learning curves for gridworld (Fig 12). As in the main results, performance is evaluated from query states that were seen in-context and at least 6 steps away from the goal. If no eligible query states meet the selection criteria, return is recorded as zero. We note that in-context learning can still be unstable at times. In part, this may be because the model is sometimes tested on states it has not experienced. Thus, it is more difficult to navigate into previously experienced territory to find the goal. We also suspect that improvements in the training procedure or architecture that we have not explored could also produce a more performant model.
@@ Appendix E Additional Behavioral Results in Gridworld and Tree Mazes
Figure 13: Additional learning results in gridworld task. A. As in Fig 2F, but for gridworld task. We also further divide the query states into states seen only before any reward experience (Pre-˘), states seen only after all reward experiences (Post-˘), states seen before and after reward experiences (Pre & Post), and states that were never seen in-context (novel). B. Depictions of the two tests for shortcut paths we use (left and right). Each test set has a fixed start location, goal location, and in-context experienced path (yellow dashed line). For each test set, we simulate 100 environments with different sensory encodings. Blue line shows an example successful shortcut path taken by the model. See methods description for more details. C. Success rate of taking the optimal, shortcut path for the two test sets in (B), across 100 sample environments. Error bars show $95\%$ confidence interval. Note that chance level in both tests is $0.02$ (generously excluding the stay transition from consideration).
@@
Appendix F Q-Learning Simulations
@@
There are also a few additional training details for the DQN. We randomly reinitialize the network weights at each context length before we run the training procedure. This is because we find that resetting the weights works empirically better than initializing with the weights from the previous context length the model was trained on (this is reasonable, as the latter induces a continual learning problem). To maintain as many parallels to standard methods as possible, we also adopt a double deep Q learning framework (van2016deep). We use a target network that is updated every $10$ epochs. We don’t think this detail is critical (and empirically the use of a target network here doesn’t seem to impact performance) as the learning problem in our setting is fully stationary.
@@
Appendix G Additional in-context representation learning results
@@
Appendix H Kernel Alignment
@@
Appendix I Analyzing Cross-Context Learning in Gridworld
@@
Appendix J Linear Decoder Setup
@@
To fit a linear decoder, we use ridge regression. We standardize features to 0 mean and unit variance. The regularization strength $\alpha$ was selected through 5-fold cross-validation using a grid search over regularization strengths from $10^{0},10^{4}]$, with 10 logarithmically-spaced values. Cross-validation was performed with shuffled splits. For each $\alpha$, we computed the mean $R^{2}$ score across all validation folds and selected the $\alpha$ that maximized this cross-validation performance. The final decoder for each layer was fit on the complete training set using the $\alpha$ found previously. Model performance was evaluated on the held-out test set.
For circular variables such as angles, we cannot directly apply standard regression since the circular nature of the data violates the assumptions of linear models (e.g., an angle of $\pi$ and $\pi$ represent the same direction but appear numerically distant). Instead, we decompose each target angle $\theta$ into its sine and cosine components: $\text{sin}(\theta)$ and $\text{cos}(\theta)$. We then fit two separate ridge regressors to predict these components independently, using the same cross-validation procedure described above. To obtain the final angle prediction, we convert the predicted sine and cosine values back to angles using the arctangent function: $)\hat{\theta}=\text{arctan2}(sin(\hat{\theta}),cos(\hat{\theta}$)).
@@
Appendix K Gradient Attribution Method
To get gradient attributions, we use integrated gradients (sundararajan2017axiomatic). As a reminder, the model output is a vector defining weights over actions. We calculate the gradient of the model’s output for the optimal action with respect to input tokens. We define the baseline inputs as the original context memory dataset $\mathcal{D}$ but with actions ablated (that is, $a=\textbf{0}$). We integrate over $20$ steps.
Appendix L Tests for Model-Free Reasoning
@@
Specifically, we assess whether the model encodes value estimates $V^{*}(s)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R_{t+1}\mid S_{0}=s\right]$ for its current state $s$ under an optimal policy. We trained linear decoders on query token representations to predict $V(s)$, but overall did not find evidence that a value gradient could be extracted from model representations (see Apps. K and L). In gridworld, $V(s)$ can be decoded with high accuracy (Appendix L). However, decoding is more accurate when $V(s)$ is defined in terms of Euclidean distance to the goal, rather than over the true 4-dimensional action space (Appendix L). This suggests that the model encodes spatial structure rather than true value gradients—its apparent $V(s)$ reflects geometric regularities, not action-contingent reward prediction. In tree mazes, decoded value estimates are localized: $V(s)$ is only reliable within 2–3 steps of the reward (Appendix L). This narrow value gradient is insufficient to guide behavior over the full task horizon.
@@
Appendix M Tests for Model-Based Reasoning
@@
We next probe for signatures of model-based reasoning. That is, we look for evidence that the model utilizes path planning to choose the correct action from the query state. This is connected to questions of state tracking (li2025language) and understanding how models simulate successive transitions between states. li2025language propose different ways that transformer models can do this path planning, from forward rollouts to more sophisticated, mergesort-like algorithms. Each of these algorithms require simulating transitions through intermediate states between query and goal. Thus, to test for the presence of path planning, we look for evidence that information about intermediate states are utilized at decision time. Specifically, we isolate decision time as computations conducted in the query token stream.
@@
Our conclusion contrasts with prior work that analyzed the behavior of meta-learned RL agents and suggested that they implement a form of model-based planning (wang2016learning; ritter2018been). It is possible that meta-learning can discover more typical model-based strategies in other settings, and these discrepancies demonstrate that the mechanistic perspective taken here can provide insights into understanding meta-learned algorithms.
Appendix N Gridworld Mechanisms
@@
In Figure 22, we show additional decoding results for gridworld. In the main text, we discussed how XY location and angle to goal can be decoded clearly from the query token stream of the model. We also find that these variables can be decoded from the context tokens. We find this interesting as it connects to our findings in tree mazes where memory tokens contain not just information about the original event (i.e., transition), but also additionally computed features.
In results section 3.5-Gridworld, we gave brief descriptions of a few analyses we ran. Here we will give more details on these analysis. First, we discussed how we showed that model performance relies on attending to tokens near the query and goal states in layer 2 (Fig. 6D). We do this by masking attention from the query token to individual context tokens, following the ablation procedure described in Section 3.4. We find that layers 1 and 3 are robust to these ablations, but performance degrades in layer 2 when attention to tokens near the query or goal is removed (Fig. 6D).
We also discussed how the attention patterns between context memory tokens shift from localized to distributed across model layers (Fig. 6), suggesting that the model first stitches transitions locally before constructing global structure. To arrive at this conclusion, we analyze the spatial locality of attention between context-memory tokens to test whether transitions are integrated locally or globally. Specifically, we plot attention strength as a function of spatial distance between token pairs (Fig. 6E). We restrict this analysis to the first two layers, since context-to-context attention in the final layer does not influence the policy output.
Appendix O Tree Maze Mechanisms
@@
In Figure 23, we show additional results for our tree maze analysis. In the tree maze task, the optimal action is often to transition to the parent node (specifically, this is true in all but the $6$ states that comprise the L-R path). This bias is reflected in the model. Without reward information, the model defaults to transition towards the parent node unless it is at the root node (Fig 23AB). Thus, we believe the model takes its default action unless it accumulates enough evidence through its layer computations to do otherwise. As discussed in the main text, we think this is done by tagging context tokens on the L-R path.
We also give more information here about analyses that we briefly described in results section 3.5-Tree Maze. We discussed a hypothesis where, at decision time, the model tests if there are context-memory tokens that contain the query state and are tagged as being on the L-R path. If so, then the correct left/right action can be inferred from the same tagged tokens (in particular since inverse actions are also encoded).
We find further evidence for this strategy by re-doing our sufficiency analysis from Fig 7C with three additional restrictions: (1) tokens must be on the L-R path in the direction to the goal, (2) tokens must be on the L-R path in the direction away from the goal, or (3) tokens are not on the L-R path at all. We find that in the first two cases the model output is unaffected, but in the third case the output is negatively impacted ((Fig 23C-F)). As long as the query token in the last layer can attend to context tokens that (1) themselves contain the query token and (2) are on the L-R path going towards or away from goal, the model output is unaffected (Fig 23C-F). The results of these perturbations are consistent with our hypothesis that decision-making in the model relies on identifying if the query state is on the L-R path via intermediate computations stored in context-memory tokens.